solve these three problems.
y= -3/2(-2)+1
y= -3/2(4)+1
y= -3/2(6)+1

Answer by Alan3354(69443)

solve these three problems.
y= -3/2(-2)+1 = 3 + 1 = 4
---------------
y= -3/2(4)+1
Multiply 4 by -3/2 and add 1.
----------------
y= -3/2(6)+1